OpenStudy (anonymous):

solve a system of equation by substitution -5x+6y=-9 -2x+y=-5

OpenStudy (anonymous):

(-2) -5x+6y=-9 (-2) (5) -2x+y=-5 (5) 10x-12y=18 -10x+5y=-25 (-7y=-7)/-7= y=1 Afterwards, take any one of the equations above and substitute y for 1 and get your x.

OpenStudy (anonymous):

thanks